### v1

@@ C.1 Descriptive Statistics
(3,546 $\times$ 10,467) (2,572 $\times$ 4,003) (3,244 $\times$ 6,464) (2,223 $\times$ 2,548)
@@
A – User–Day Level (N = 186,496)
   Output 0.0122 0.1099 0.0072 0.0847 0.0152 0.1223 0.0189 0.1361
   Output (Shen, 2023) 0.0152 0.1224 0.0101 0.1002 0.0182 0.1338 0.0237 0.1521
   Commit 0.0095 0.0968 0.0053 0.0726 0.0119 0.1086 0.0144 0.1192
   Issue closed 0.0035 0.0594 0.0021 0.0455 0.0044 0.0663 0.0052 0.0722
   Bug fix ratio 0.0010 0.0305 0.0006 0.0250 0.0012 0.0333 0.0015 0.0378
   PR merged 0.0048 0.0690 0.0023 0.0476 0.0063 0.0789 0.0078 0.0878
   PR merge ratio 0.4620 6.7223 0.2235 4.7031 0.6038 7.6715 0.7471 8.5227
   Files edited per PR 0.0095 0.1534 0.0055 0.1239 0.0119 0.1685 0.0149 0.1822
   Lines added per PR 0.0211 0.3244 0.0124 0.2588 0.0263 0.3577 0.0331 0.3938
B – User Level (N = 10,467)
@@
   User creation year 2015.39 3.51 2009 2015 2023
@@
   Followers 53.83 343.70 0 11 17421
   Following 37.16 325.79 0 11 28300
   Repositories 40.66 78.77 0 23 3900
   Total events 3.73 10.59 0 1 138
@@
  • Panel A presents descriptive statistics for the baseline sample period March 27–30 (Pre) and April 3–6 (Post). The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a(n) (analytical) programming package repository. The number of unique repository $\times$ GitHub user accounts for the entire repository–user-level sample (“Overall”) and for each of the subsamples is presented in parentheses below. Panel B provides information on the individual characteristics of all GitHub user accounts in the repository–user-level sample.
C.2 Additional Results
Table C.2: Effect of ChatGPT Ban on GitHub Output — Repository-Level Analysis
Output Quantity Output Quality Task Complexity
Output Output (Shen, 2023) Commit Issue closed Bug fix ratio PR merged PR merge ratio Files edited per PR Lines added per PR
(1) (2) (3) (4) (5) (6) (7) (8) (9)
Pre Mar 27–28 – Post Apr 3–4
Overall (N = 93,248) Treated $\times$ Post -0.0011 -0.0012 -0.0018 -0.0006 0.0001 -0.0008 -0.0667 -0.0007 -0.0033
(0.0015) (0.0018) (0.0014) (0.0009) (0.0004) (0.0010) (0.0957) (0.0022) (0.0047)
Less experienced (N = 34,768) Treated $\times$ Post 0.0006 0.0012 -0.0006 0.0002 0.0003 0.0011 0.1042 0.0050 0.0096
(0.0020) (0.0025) (0.0018) (0.0010) (0.0006) (0.0012) (0.1128) (0.0029) (0.0065)
Experienced (N = 58,480) Treated $\times$ Post -0.0020 -0.0026 -0.0023 -0.0011 $0$0.0019 -0.1649 -0.0044 -0.0114
(0.0022) (0.0024) (0.0019) (0.0013) (0.0006) (0.0015) (0.1416) (0.0031) (0.0065)
Pkg. contributor (N = 23,668) Treated $\times$ Post -0.0039 -0.0053 -0.0037 -0.0017 -0.0026 -0.0003 -0.0078 -0.0011 -0.0025
(0.0039) (0.0045) (0.0035) (0.0023) (0.0012) (0.0025) (0.2399) (0.0050) (0.0103)
Pre Mar 27–30 – Post Apr 3–6
Overall (N = 186,496) Treated $\times$ Post -0.0014 -0.0010 -0.0013 -0.0007 0.0001 -0.0002 -0.0190 -0.0007 -0.0022
(0.0012) (0.0014) (0.0011) (0.0007) (0.0003) (0.0007) (0.0704) (0.0017) (0.0035)
Less experienced (N = 69,536) Treated $\times$ Post 0.0019 0.0026 0.0008 0.0016 0.0007 0.0014 0.1319 0.0038 0.0085
(0.0016) (0.0019) (0.0014) (0.0008) (0.0004) (0.0008) (0.0828) (0.0025) (0.0049)
Experienced (N = 116,960) Treated $\times$ Post -0.0035 -0.0032 -0.0025 -0.0022 -0.0002 -0.0012 -0.1098 -0.0034 -0.0089
(0.0017) (0.0019) (0.0016) (0.0010) (0.0004) (0.0011) (0.1042) (0.0023) (0.0047)
Pkg. contributor (N = 47,336) Treated $\times$ Post -0.0051 -0.0031 -0.0031 -0.0023 -0.0011 0.0002 0.0267 -0.0031 -0.0052
(0.0031) (0.0035) (0.0029) (0.0017) (0.0008) (0.0018) (0.1735) (0.0038) (0.0079)
@@ C.2 Additional Results
  • All specifications include repository $\times$ user fixed effects. The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a(n) (analytical) programming package repository. The number of observations is in parentheses after each sample definition. Robust standard errors in parentheses are clustered at the repository $\times$ user level.
Table C.3: Repository-Level Analysis — March 27–28 (Pre) – April 3–4 (Post)
Output Quantity Output Quality Task Complexity
Output Output (Shen, 2023) Commit Issue closed Bug fix ratio PR merged PR merge ratio Files edited per PR Lines added per PR
(1) (2) (3) (4) (5) (6) (7) (8) (9)
Treated $\times$ Post -0.0011 -0.0012 -0.0018 -0.0006 0.0001 -0.0008 -0.0667 -0.0007 -0.0033
(0.0015) (0.0018) (0.0014) (0.0009) (0.0004) (0.0010) (0.0957) (0.0022) (0.0047)
Post 0.0008 0.0011 0.0009 0.0001 0.0001 0.0007 0.0631 0.0021 0.0041
(0.0007) (0.0008) (0.0006) (0.0004) (0.0002) (0.0005) (0.0447) (0.0012) (0.0024)
Overall (N = 93,248) Dep. var. mean 0.0124 0.0155 0.0095 0.0036 0.0010 0.0048 0.4618 0.0097 0.0212
Treated $\times$ Post 0.0006 0.0012 -0.0006 0.0002 0.0003 0.0011 0.1042 0.0050 0.0096
(0.0020) (0.0025) (0.0018) (0.0010) (0.0006) (0.0012) (0.1128) (0.0029) (0.0065)
Post -0.0012 -0.0012 -0.0010 0.0000 -0.0003 -0.0009 -0.0864 -0.0010 -0.0034
(0.0008) (0.0010) (0.0007) (0.0005) (0.0003) (0.0005) (0.0515) (0.0012) (0.0026)
Less experienced (N = 34,768) Dep. var. mean 0.0075 0.0106 0.0056 0.0020 0.0007 0.0024 0.2373 0.0053 0.0121
Treated $\times$ Post -0.0020 -0.0026 -0.0023 -0.0011 $0$0.0019 -0.1649 -0.0044 -0.0114
(0.0022) (0.0024) (0.0019) (0.0013) (0.0006) (0.0015) (0.1416) (0.0031) (0.0065)
Post 0.0020 0.0023 0.0019 0.0002 0.0004 0.0016 0.1457 0.0037 0.0082
(0.0010) (0.0011) (0.0008) (0.0006) (0.0003) (0.0007) (0.0633) (0.0017) (0.0034)
Experienced (N = 58,480) Dep. var. mean 0.0153 0.0184 0.0118 0.0046 0.0011 0.0062 0.5953 0.0123 0.0267
Treated $\times$ Post -0.0039 -0.0053 -0.0037 -0.0017 -0.0026 -0.0003 -0.0078 -0.0011 -0.0025
(0.0039) (0.0045) (0.0035) (0.0023) (0.0012) (0.0025) (0.2399) (0.0050) (0.0103)
Post 0.0039 0.0024 0.0024 0.0003 0.0006 0.0016 0.1394 0.0058 0.0118
(0.0016) (0.0020) (0.0014) (0.0010) (0.0005) (0.0011) (0.1053) (0.0025) (0.0052)
Pkg. contributor (N = 23,668) Dep. var. mean 0.0192 0.0243 0.0142 0.0055 0.0014 0.0077 0.7437 0.0147 0.0326
@@
  • All specifications include repository $\times$ user fixed effects. The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a(n) (analytical) programming package repository. The number of observations is in parentheses after each sample definition. Robust standard errors in parentheses are clustered at the repository $\times$ user level.
Table C.4: Repository-Level Analysis — March 27–30 (Pre) – April 3–6 (Post)
Output Quantity Output Quality Task Complexity
Output Output (Shen, 2023) Commit Issue closed Bug fix ratio PR merged PR merge ratio Files edited per PR Lines added per PR
(1) (2) (3) (4) (5) (6) (7) (8) (9)
Treated $\times$ Post -0.0014 -0.0010 -0.0013 -0.0007 0.0001 -0.0002 -0.0190 -0.0007 -0.0022
(0.0012) (0.0014) (0.0011) (0.0007) (0.0003) (0.0007) (0.0704) (0.0017) (0.0035)
Post 0.0004 0.0004 0.0003 0.0002 0.0002 0.0003 0.0322 0.0021 0.0036
(0.0005) (0.0006) (0.0005) (0.0003) (0.0002) (0.0003) (0.0315) (0.0008) (0.0016)
Overall (N = 186,496) Dep. var. mean 0.0122 0.0152 0.0095 0.0035 0.0010 0.0048 0.4620 0.0095 0.0211
Treated $\times$ Post 0.0019 0.0026 0.0008 0.0016 0.0007 0.0014 0.1319 0.0038 0.0085
(0.0016) (0.0019) (0.0014) (0.0008) (0.0004) (0.0008) (0.0828) (0.0025) (0.0049)
Post -0.0013 -0.0017 -0.0011 -0.0004 -0.0002 -0.0008 -0.0763 -0.0005 -0.0026
(0.0007) (0.0008) (0.0006) (0.0004) (0.0002) (0.0004) (0.0355) (0.0009) (0.0020)
Less experienced (N = 69,536) Dep. var. mean 0.0072 0.0101 0.0053 0.0021 0.0006 0.0023 0.2235 0.0055 0.0124
Treated $\times$ Post -0.0035 -0.0032 -0.0025 -0.0022 -0.0002 -0.0012 -0.1098 -0.0034 -0.0089
(0.0017) (0.0019) (0.0016) (0.0010) (0.0004) (0.0011) (0.1042) (0.0023) (0.0047)
Post 0.0014 0.0016 0.0011 0.0005 0.0005 0.0010 0.0921 0.0035 0.0070
(0.0007) (0.0008) (0.0006) (0.0004) (0.0002) (0.0005) (0.0447) (0.0011) (0.0023)
Experienced (N = 116,960) Dep. var. mean 0.0152 0.0182 0.0119 0.0044 0.0012 0.0063 0.6038 0.0119 0.0263
Treated $\times$ Post -0.0051 -0.0031 -0.0031 -0.0023 -0.0011 0.0002 0.0267 -0.0031 -0.0052
(0.0031) (0.0035) (0.0029) (0.0017) (0.0008) (0.0018) (0.1735) (0.0038) (0.0079)
Post 0.0027 0.0008 0.0008 0.0005 0.0004 0.0008 0.0773 0.0054 0.0106
(0.0013) (0.0015) (0.0011) (0.0007) (0.0004) (0.0009) (0.0812) (0.0018) (0.0040)
Pkg. contributor (N = 47,336) Dep. var. mean 0.0189 0.0237 0.0144 0.0052 0.0015 0.0078 0.7471 0.0149 0.0331
  • Notes:
  • All specifications include repository $\times$ user fixed effects. The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a(n) (analytical) programming package repository. The number of observations is in parentheses after each sample definition. Robust standard errors in parentheses are clustered at the repository $\times$ user level.
Figure C.1: Repository-Level Analysis
Notes: Panel A displays event-study estimates across outcomes for “less experienced” GitHub user accounts (created after or in 2017). Panel B presents event-study estimates for (i) “experienced” GitHub user accounts (created before 2017) and (ii) and accounts that are the owner and/or contributor to a(n) (analytical) programming package repository (“Pkg. contributor”). The sample period spans March 27–30 (Pre) and April 3–6 (Post). All specifications include user $\times$ repository, time, and day-of-the-week fixed effects; 95% (90%) confidence intervals for robust standard errors clustered at the user $\times$ repository level are depicted in light (dark) grey.
Appendix D User Adaption to the ChatGPT Ban
@@ D.1 Data
Figure D.1: Panel Structure of Google Trends and TOR Data
Notes: The panel structure of the datasets used in Section D—i.e., the (i) Google trends and the (ii) TOR (“standard” and bridge relay) user datasets—are displayed. Workday dates during the sample period are presented on the x-axis, and ISO 3166-1 alpha-2 country codes are presented on the y-axis.
D.2 Results
@@
The top panel in Figure D.2 presents the effect of the ChatGPT ban on the number of Google queries on the topic of VPNs. We observe a significant increase in the share of web searches on this relative to other topics in Italy on the first working day after the ban that slowly vanishes over the next three days. The estimated effect on April 3 is sizeable: the share of searches on VPNs increases by 52.2 percentage points. On average, the share of queries on VPNs was 20.6 percentage points higher in Italy over the workweek. The observed pattern is consistent with Italian users looking for ways to access ChatGPT even after the ban and succeeding after some initial search costs. Our estimates might, however, present only stated preferences.
Figure D.2: Effect of ChatGPT Ban on Ban Circumvention Technology
@@
To investigate whether the ban actually led to behavioral changes among Italian users, we look at an alternative outcome: the log number of TOR users. The results for TOR relay and TOR bridge users are presented in Panels B and C of Figure D.2, respectively. While the number of TOR relay users shows only a minor increase in the days after the ban, the average treatment effect on the number of Italian TOR bridge users is positive and significant on the first workday after the ban. Usage of TOR bridges remained elevated for the entire workweek, with an increase in user numbers—on average—of approximately 9.4 percentage points. This pattern is in line with users resorting to bridge over “standard” relays to minimize the chance of their being denied access to ChatGPT since the former are more difficult for firewalls to identify. [^28] For a discussion on denial of ChatGPT access, see the following OpenAI forum discussion: https://community.openai.com/t/access-denied-error-1020/38758/23.

### v2

@@ Appendix D GitHub Repository–User-Level Data
(N = 4,627 $\times$ 11,938) (N = 3,315 $\times$ 4,566) (N = 4,294 $\times$ 7,372) (N = 3,006 $\times$ 2,902)
@@
A - Repo-User-Day–Level (N = 239,912)
     Output 0.0116 0.1069 0.0080 0.0889 0.0136 0.1158 0.0171 0.1296
     Output (Shen, 2023) 0.0108 0.1036 0.0075 0.0860 0.0128 0.1122 0.0163 0.1264
     Output (Holub and Thies, 2023) 0.0279 0.1647 0.0192 0.1373 0.0327 0.1780 0.0405 0.1971
     Issue closed 0.0024 0.0492 0.0014 0.0376 0.0030 0.0546 0.0034 0.0580
     PR merge ratio 0.0030 0.0542 0.0013 0.0358 0.0039 0.0621 0.0047 0.0679
     PR opened 0.0054 0.0734 0.0041 0.0641 0.0061 0.0780 0.0080 0.0890
     Avg. lines added per opened PR 0.0170 0.2762 0.0134 0.2494 0.0190 0.2902 0.0238 0.3190
     Avg. lines added per merged PR 0.0104 0.2167 0.0047 0.1483 0.0135 0.2469 0.0153 0.2541
     Easy issue closed 0.0001 0.0094 0.0000 0.0034 0.0001 0.0114 0.0001 0.0089
     Interactive Activity 0.0193 0.1377 0.0127 0.1120 0.0230 0.1500 0.0285 0.1663
B - User–Level (N = 11,938)
@@
     User creation year 2015.37 3.51 2009 2015 2023
@@
     Followers 49.65 300.24 0 12 17421
     Following 36.54 313.69 0 11 28300
     Repositories 40.40 78.31 0 23 3900
     Total events 2.89 7.47 0 1 98
@@
  • Panel A presents descriptive statistics for the baseline sample period Pre 27-30.03 – Post 03-06.04. The “Less experienced” sample includes all GitHub user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of unique GitHub user accounts for the entire repository $\times$ user sample (“Overall”) and each of the subsamples is presented in parentheses below. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Panel B provides information on the individual characteristics of all GitHub user accounts in the repository $\times$ user sample.
(1) (2) (3) (4) (5)
A: Output Quantity and Quality
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Overall Treated $\times$ Post -0.0010 -0.0007 0.0029* -0.0007 -0.0002
(N = 239,912) (0.0010) (0.0010) (0.0016) (0.0005) (0.0005)
Dep. var. mean 0.0116 0.0108 0.0279 0.0024 0.0030
Less experienced Treated $\times$ Post 0.0019 0.0014 0.0047** 0.0007 0.0014***
(N = 86,200) (0.0015) (0.0014) (0.0022) (0.0006) (0.0005)
Dep. var. mean 0.0080 0.0075 0.0192 0.0014 0.0013
Experienced Treated $\times$ Post -0.0027** -0.0019 0.0018 -0.0016** -0.0011
(N = 153,712) (0.0014) (0.0013) (0.0022) (0.0007) (0.0007)
Dep. var. mean 0.0136 0.0128 0.0327 0.0030 0.0039
Pkg. contributor Treated $\times$ Post -0.0036 -0.0021 0.0011 -0.0025** 0.0001
(N = 63,064) (0.0025) (0.0024) (0.0036) (0.0012) (0.0011)
Dep. var. mean 0.0171 0.0163 0.0405 0.0034 0.0047
B: Task Choice and Complexity
PR opened Avg. lines added per opened PR Avg. lines added per merged PR Easy issue closed Interactive Activity
Overall Treated $\times$ Post 0.0008 0.0022 -0.0008 0.0001 0.0027**
(N = 239,912) (0.0007) (0.0027) (0.0020) (0.0001) (0.0013)
Dep. var. mean 0.0054 0.0170 0.0104 0.0001 0.0193
Less experienced Treated $\times$ Post 0.0019* 0.0059 0.0073*** 0.0000 0.0028
(N = 86,200) (0.0011) (0.0044) (0.0024) (0.0000) (0.0018)
Dep. var. mean 0.0041 0.0134 0.0047 0.0000 0.0127
Experienced Treated $\times$ Post 0.0001 0.0002 -0.0058** 0.0001 0.0026
(N = 153,712) (0.0010) (0.0034) (0.0029) (0.0002) (0.0018)
Dep. var. mean 0.0061 0.0190 0.0135 0.0001 0.0230
Pkg. contributor Treated $\times$ Post 0.0003 -0.0030 -0.0015 -0.0002 0.0031
(N = 63,064) (0.0016) (0.0058) (0.0043) (0.0002) (0.0031)
Dep. var. mean 0.0080 0.0238 0.0153 0.0001 0.0285
@@
  • All specifications include repository $\times$ user fixed effects and day-of-the-week–fixed effects. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the repository $times$ user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
(1) (2) (3) (4) (5)
Output Output (Shen, 2023) Output (Holub & Thies, 2023 Issue closed PR merge ratio
Treated $\times$ Post -0.0042** -0.0022 -0.0007 -0.0021** -0.0011
(0.0021) (0.0021) (0.0040) (0.0009) (0.0007)
Overall (N = 33,512) Dep. var. mean 0.0078 0.0073 0.0227 0.0016 0.0016
Treated $\times$ Post -0.0004 0.0017 0.0027 -0.0019* 0.0004
(0.0029) (0.0028) (0.0057) (0.0010) (0.0007)
Less experienced (N = 12,784) Dep. var. mean 0.0055 0.0052 0.0173 0.0007 0.0003
Treated $\times$ Post -0.0069** -0.0048* -0.0031 -0.0023* -0.0021*
(0.0029) (0.0029) (0.0056) (0.0013) (0.0011)
Experienced (N = 20,728) Dep. var. mean 0.0093 0.0086 0.0261 0.0022 0.0023
Treated $\times$ Post -0.0060 -0.0024 0.0057 -0.0037** -0.0025
(0.0042) (0.0042) (0.0082) (0.0019) (0.0017)
Pkg. contributor (N = 11,392) Dep. var. mean 0.0121 0.0123 0.0364 0.0011 0.0020
@@
  • All specifications include repository $\times$ user fixed effects and day-of-the-week–fixed effects. The “Less experienced” sample includes all Github user accounts created after or in 2017 (median), while the “Experienced” sample comprises all GitHub user accounts created before 2017. The “Pkg. contributor” sample comprises all GitHub user accounts that are the owner and/or contributor to a (analytical) programming package repository. The number of observations is depicted in parentheses after each sample definition. A log plus one transformation is applied to Avg. lines added per PR (opened or merged). Robust standard errors in parentheses are clustered on the repository $times$ user-level: * $p<0.1$, ** $p<0.05$, *** $p<0.01$.
Appendix E User Adaption to the ChatGPT Ban
@@ E.1 Data
Notes: The panel structure of the datasets used in Section LABEL:sec:google-tor—i.e., the (i) Google trends and the (ii) TOR (“standard” and bridge relay) user datasets—are displayed. Workday dates during the sample period are presented on the x-axis, and ISO 3166-1 alpha-2 country codes are presented on the y-axis.
E.2 Results
@@
The top panel in Figure LABEL:fig:ban-circumvention presents the effect of the ChatGPT ban on the number of Google queries on the topic of VPNs. We observe a significant increase in the share of web searches on this relative to other topics in Italy on the first working day after the ban that slowly vanishes over the next three days. The estimated effect on April 3 is sizeable: the share of searches on VPNs increases by 52.2 percentage points. On average, the share of queries on VPNs was 20.6 percentage points higher in Italy over the workweek. The observed pattern is consistent with Italian users looking for ways to access ChatGPT even after the ban and succeeding after some initial search costs. Our estimates might, however, present only stated preferences.
@@
To investigate whether the ban actually led to behavioral changes among Italian users, we look at an alternative outcome: the log number of TOR users. The results for TOR relay and TOR bridge users are presented in Panels B and C of Figure LABEL:fig:ban-circumvention, respectively. While the number of TOR relay users shows only a minor increase in the days after the ban, the average treatment effect on the number of Italian TOR bridge users is positive and significant on the first workday after the ban. Usage of TOR bridges remained elevated for the entire workweek, with an increase in user numbers—on average—of approximately 9.4 percentage points. This pattern is in line with users resorting to bridge over “standard” relays to minimize the chance of their being denied access to ChatGPT since the former are more difficult for firewalls to identify. [^36] For a discussion on denial of ChatGPT access, see the following OpenAI forum discussion: https://community.openai.com/t/access-denied-error-1020/38758/23.